INFRARED SPECTRA OF He–CS_{2}, Ne–CS_{2}, AND Ar–CS_{2}

F. MIVEHVAR, \underline{J. NOROOZ OLIAEE}, N. MOAZZEN-AHMADI, Department of Physics and Astronomy, University of Calgary, Calgary, AB T2N 1N4, Canada.

Infrared spectra of weakly bound Rg–CS_{2} (Rg = He, Ne, and Ar) clusters formed in a pulsed supersonic slit-jet expansion have been recorded by exciting the CS_{2} ν_{3} fundamental band (∼ 1535 cm^{−1}) using a tuneable diode laser. Spectra were well fitted to a conventional semi-rigid asymmetric rotor Hamiltonian. The He–CS_{2} spectrum was assigned to an a-type band, while spectra of Ne–CS_{2} and Ar–CS_{2} were well described by b-type bands, indicating a/b axis switching in transition from the He–CS_{2} complex to the Ne–CS_{2} and Ar–CS_{2} complexes. The results show that the complexes have vibrationally averged T-shaped structures. The determined structural parameters along with the observed vibrational shifts are R = 3.81, 3.57 and 3.71 Å, θ = 80.0, 86.9 and 86.4^{∘} and Δν = 0.171, 0.181 and 0.067 cm^{−1} for He–CS_{2}, Ne–CS_{2} and Ar–CS_{2}, respectively. Here, R is the distance between the rare gas and the carbon atom, θ is the the angle between R and and the CS_{2} axis and Δν is the vibrational shift with respect to the free CS_{2} monomer.